### v1

# AlignXIE: Improving Multilingual Information Extraction by Cross-Lingual Alignment

Abstract
Empirical evidence suggests that LLMs exhibit spontaneous cross-lingual alignment. Our findings suggest that although LLMs also demonstrate promising cross-lingual alignment in Information Extraction, there remains significant imbalance across languages, revealing an underlying deficiency in the IE alignment. To address this issue, we propose AlignXIE , a powerful code-based LLM that significantly enhances cross-lingual IE alignment through two strategies. Firstly, AlignXIE formulates IE across different languages, especially non-English ones, as code generation tasks, standardizing the representation of various schemas using Python classes to ensure consistency of the same ontology in different languages and align the schema. Secondly, it incorporates an IE cross-lingual alignment phase through a translated instance prediction task proposed in this paper to align the extraction process, utilizing ParallelNER, an IE bilingual parallel dataset with 257,190 samples, generated by our proposed LLM-based automatic pipeline for IE parallel data construction, with manual annotation to ensure quality. Ultimately, we obtain AlignXIE through multilingual IE instruction tuning. Although without training in $9$ unseen languages, AlignXIE surpasses ChatGPT by 30.17% and SoTA by 20.03%, thereby demonstrating superior cross-lingual IE capabilities. Comprehensive evaluations on 63 IE benchmarks in Chinese and English under various settings, demonstrate that AlignXIE significantly enhances cross-lingual and multilingual IE through boosting the IE alignment.
{CJK}
UTF8gkai
AlignXIE: Improving Multilingual Information Extraction by Cross-Lingual Alignment
Yuxin Zuo1thanks:   Co-first authors., Wenxuan Jiang211footnotemark: 1, Wenxuan Liu [^1] footnotemark: 1, Zixuan Li1, Long Bai1, Hanbin Wang3, Yutao Zeng 1, Xiaolong Jin1thanks:   Corresponding author., Jiafeng Guo1, Xueqi Cheng1 1Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences 2School of Software, Northeastern University 3School of Software, Peking University {zuoyuxin24s, liuwenxuan2024z, jinxiaolong}@ict.ac.cn
1 Introduction
CoNLL 2003 CoNLL 2003-zh
Flan-T5 12.71 2.13
InstructUIE 92.94 20.39
Table 1: Results on CoNLL 2003 and CoNLL 2003-zh, the latter being the constructed Chinese parallel dataset.
Multilingual Information Extraction (Multilingual IE) focuses on automatically extracting structured knowledge from unstructured or semi-structured texts across various languages, following manually designed schemas. Large Language Models (LLMs) (Brown et al., 2020; Touvron et al., 2023a, b), trained on massive multilingual corpora, have significantly advanced multilingual language processing. Previous studies (Zhang et al., 2023; Qi et al., 2023; Wang et al., 2024; Gao et al., 2024) have shown that LLMs exhibit spontaneous cross-lingual alignment to facilitate the transfer of abilities and knowledge across languages. Our findings suggest the presence of this alignment in IE, indicating a strong potential for improving the IE cross-lingual transfer. We first define IE parallel data, which is across various languages that share the same schema, sentences, and extracted instances. Previous studies proposed label projection to generate such data to alleviate the challenges of IE in low-resource Kolluru et al. (2022); Hennig et al. (2023); Rios et al. (2024). By translation and manual annotation, we constructed an IE parallel dataset in Chinese based on CoNLL 2003 Sang and Meulder (2003), to evaluate the InstructUIE Wang et al. (2023a). Table 1 shows the results. After training Flan-T5 Chung et al. (2022) on English IE to obtain InstructUIE, there was a notable enhancement in Chinese IE ($2.13$$\rightarrow$20.3920.3920.3920.39), strongly supporting the presence of IE cross-lingual alignment. However, the significant performance gaps between languages ($92.94$ vs $20.39$) indicate that the alignment remains weak. To address this issue, we propose two strategies to enhance IE cross-lingual alignment by aligning the schema and extraction process in different languages, thereby improving the cross-lingual and multilingual IE.
Firstly, we formulate IE tasks across different languages, particularly non-English languages, into a unified framework of code generation. Traditional multilingual IE faces a major challenge from cross-lingual schema differences, requiring schema alignment before knowledge sharing, which hinders cross-lingual transfer. Yang et al. found that unified schema representations in a single language can significantly enhance model generalization, and thus, we construct a unified schema representation to align cross-lingual schemas. Leveraging the language-agnostic nature of code, we use the Python classes to represent schemas across various languages, which ensures consistent representation of the same or similar ontology in different languages, thereby facilitating knowledge transfer for IE. The consistent schema allows the model to learn and efficiently share the knowledge of the same ontology across languages.
Secondly, we incorporate an IE cross-lingual alignment instruction tuning phase. We propose a task utilizing the IE parallel data, aiming at predicting target language instances to facilitate the alignment of the extraction process. Leveraging prediction tasks with parallel data to enhance alignment through instruction tuning has become a common paradigm in multilingual processing (Zhu et al., 2023; Gao et al., 2024). Figure 1 presents an example. Similar to the label projection, the challenge of the task is to ensure that the predicted instances remain semantically consistent with the source-language ones while appearing in the target language sentence. In contrast to predicting complete IE parallel data directly, which mainly focuses on the alignment of the target-language sentence that will be provided in IE, our method emphasizes the alignment of translated instances, which is the core objective of IE cross-lingual alignment.
To use high-quality IE parallel data for the phase, we propose an automatic three-stage LLM-based pipeline that implements contextual translation and multi-grained rephrasing. The pipeline achieved $99\%$ average faithfulness across $10$ languages evaluated on the label projection benchmark WikiANN Pan et al. (2017). Utilizing the pipeline, we construct a high-quality and diverse NER English-Chinese parallel dataset ParallelNER, using the GPT-4 series model supplemented with manual annotation. By distilling the IE alignment capabilities of advanced proprietary models, the IE cross-lingual alignment phase enhances cross-lingual generalization in IE tasks.
Ultimately, we obtain the AlignXIE through multilingual IE instruction tuning. We evaluate the AlignXIE on $63$ benchmarks for the NER, RE, ED, and EAE tasks in Chinese and English. AlignXIE exhibits superior cross-lingual generalization across typologically diverse languages, and surpasses ChatGPT by $30.17\%$ and SoTA by $20.03\%$. In the supervised evaluation, AlignXIE consistently ranks within the top-$2$ results across $40$ (of $42$) benchmarks, with an average improvement of $3.03$ and $2.92$ points on the English RE and ED tasks, respectively. Notably, AlignXIE achieves SoTA across all Chinese IE benchmarks under all settings, effectively demonstrating the strength of knowledge transfer from English IE through our method.
Our contributions can be summarized as:
  • The code-based multilingual IE method of AlignXIE unifies the representation of schemas across different languages, thereby boosting cross-lingual transfer.
  • AlignXIE incorporates a cross-lingual alignment phase for IE to improve cross-lingual transfer. This phase involves instruction tuning on a newly proposed translated instance prediction task.
  • We propose an LLM-based automatic IE parallel data construction pipeline. Using this pipeline, we constructed a large-scale, high-quality bilingual parallel NER dataset ParallelNER with $257,190$ samples to provide valuable resources to the community.
Figure 1: An example of instruction-tuning data used in the IE cross-lingual alignment phase. The highlighted span in the completion and the two same color highlighted spans in the instruction must ensure semantic consistency and be exactly matched, respectively.
2 Code-Based Multilingual IE
AlignXIE formulates the IE tasks across various languages using a unified code generation framework. In this section, we introduce the instruction and completion of the proposed code-based multilingual IE. The instruction consists of two parts (§2.1): (1) representation code of schema across different languages (§2.1.1); (2) task prompts for completion (§2.1.2). Subsequently, we present the completion format (§2.2).
2.1 Instruction Format
2.1.1 Schema Representation
Class Definition
The IE schema comprises multiple concepts, each of which consists of a name and a set of attributes. The three fundamental concepts of IE schemas [Entity, Relation, Event] are first defined as base classes. For each concept of a schema, a corresponding class is subsequently defined based on its name. Each concept inherits from its respective base class by default. Then, we define the attributes of each concept (such as argument roles) as input arguments of the constructor __init__(). Specially, for IE in non-English languages, we first create a mapping from the original schema to the English schema through translation or manual annotation. The mapping standardizes the representation of the same ontology across different languages by utilizing the unified Python class. For example, the entity concept {CJK}UTF8mj사람 in Korean and {CJK}UTF8gkai人物 in Chinese are both mapped to the class PER(Entity), ensuring consistent semantic representation across linguistic boundaries to align the schema in different languages.
Class Comments
We utilize class comments to provide clear definitions of concepts similar to GoLLIE Sainz et al. (2024) and KnowCoder Li et al. (2024). Class comments comprise two components: 1) a description that explains the corresponding concept; 2) examples contain instances corresponding to the concept. For the description part, we follow the annotation guidelines of datasets by default. We propose an LLM-based approach to generate descriptions for datasets where guidelines are unavailable. To enable the LLM to better understand the concepts and the annotation style of the dataset, we first sample $10$ instances to prompt the LLM to generate an initial description. Subsequently, to avoid providing excessive examples at once to hinder effective induction, we sample $20$ instances for the LLM to iteratively refine the description based on each instance. Specifically, for each instance, if it cannot be categorized to the corresponding concept according to the description, the LLM will adjust the inappropriate parts of the current description accordingly. We use the GPT-4o-2024-08-06 as the LLM. The prompts are shown in Appendix C. For the examples part, we sample up to $10$ instances for each concept by frequency.
2.1.2 Task Prompt
The task prompt contains a docstring that defines the task and incorporates a string variable sentence in Python, which contains the text to be extracted. Then, the task prompt is concatenated to the schema, resulting in the instruction.
2.2 Completion Format
The completion represents the final output generated by the model, which begins after the assignment operation (i.e., results =). Each completion is a list, where each element is an instantiated object, representing the extracted structured knowledge instance. Additionally, we utilize the exec() function in Python to execute the instruction and completion of all code-based data in this work to ensure quality. We also use this function to execute the generated code to extract the results.
3 Training Framework
AlignXIE adopts a two-phase training framework based on instruction tuning. The first phase is the IE cross-lingual alignment phase, where we propose the task and data (§3.1). Subsequently, we conduct the multilingual IE instruction tuning to obtain AlignXIE  (§3.2). Appendix D shows examples of instruction tuning data of the two phases.
3.1 IE Cross-Lingual Alignment Phase
First, we introduce the instruction and completion of the translated spans prediction task (§3.1.1). To obtain high-quality IE parallel data for this training phase, we propose a three-stage automatic pipeline for constructing IE parallel data (§3.1.2) and create a dataset ParallelNER (§3.1.3).
3.1.1 Translated Instances Prediction Task
The translated instances prediction task is the training objective of this phase, aiming to perform IE in the target language. The instruction starts with a task description, prompting the LLM to predict translated instances by concatenating the input-output pairs of source-language IE data with the input of target-language IE data to form the instruction. The completion contains the output of the target language data. Figure 1 shows an example.
Initialize : Source Sentence $s^{src}$
Source Spans $I^{src}$
Output : Target Sentence $s^{tgt}$
Target Spans $I^{tgt}$
Function Translate($s^{src},I^{src}$):
       $prompt=$ TemplateT ($s^{src},I^{src}$) ;
       // Generate prompt for Contextual Translation.
       $I^{tgt},s^{tgt}=\textnormal{{LLM}}(prompt)$;
       return $I^{tgt},s^{tgt}$ ;
Function SpanRephrase($I^{tgt},I^{src},s^{tgt}$):
       $Flag=False$ ;
       foreach $(i^{tgt}_{m},i^{src}_{m})$ in $(I^{tgt},I^{src})$ do
             if $i^{tgt}_{m}$ not in $s^{tgt}$ then
                   $prompt=$ TemplateI ($s^{tgt},i^{src}_{m},i^{tgt}_{m}$);
                   // Generate prompt for Span Rephrase. $i^{cor}_{m}=$ LLM ($prompt$) ;
                  if $i^{cor}_{m}$ in $s^{tgt}$ then
                         $I^{tgt}.update(i^{tgt}_{m},i^{cor}_{m})$ ;
                  else
                         $Flag=$ True ;
      return $I^{tgt},Flag$ ;
Function SentenceRephrase($I^{tgt},s^{src}$):
       $prompt=$ TemplateS ($I^{tgt},s^{src}$) ; // Generate prompt for Sentence Rephrase. $s^{tgt}=$ LLM ($prompt$); return $s^{tgt}$ ;
Function Main($s^{src},I^{src}$):
       $I^{tgt},s^{tgt}=$ Translate($s^{src},I^{src}$);
       $I^{tgt},Flag=$ SpanRephrase($I^{tgt},I^{src},s^{tgt}$);
       if Flag == True then
             $s^{tgt}=$ SentenceRephrase($I^{tgt},s^{src}$);
      return $s^{tgt},I^{tgt}$
Algorithm 1 Span Translation
3.1.2 IE Parallel Data Construction Pipeline
The construction of IE parallel data involves two critical components: span translation and schema translation. Spans refer to the desirable spans annotated in the text of the IE task Kripke and Munitz (1971); Chen and Yuille (2004). Our pipeline mainly performs span translation, followed by a post-processing step that translates the schema and associates it with the corresponding spans Onyshkevych (1994); Milward and Thomas (2000); Lu et al. (2022). Schema translation can be accomplished using LLMs and manual annotation. The primary challenge lies in the span translation, which involves two key issues: 1) Inaccurate translation: Either the spans or the sentences are inaccurately translated; 2) Missing Spans: The translated spans may be missing from the translated sentence. To address these issues, we propose a three-stage pipeline implementing contextual translation and multi-grained rephrasing, which is capable of generating parallel data for any IE task across any languages, as shown in Algorithm 1. The detailed prompts are shown in Appendix C.
Stage 1: Contextual Translation
We conduct the contextual translation first to ensure the semantic consistency of the translated sentence and spans. Given the sentence $s^{src}$ and spans list $I^{src}=\{i^{src}_{1},i^{src}_{2},...,i^{src}_{n}\}$ in the source language, the aim of this stage is to obtain the initial sentence $s^{tgt}$ and spans list $I^{tgt}=\{i^{tgt}_{1},i^{tgt}_{2},...,i^{tgt}_{n}\}$ in the target language. We use a predefined template to format the sentence $s^{src}$ with spans $I^{src}$ as the prompt for the LLM to translate jointly.
Stage 2: Span Rephrase
The latter two stages address the challenge of missing spans through rephrasing. This stage focuses on rephrasing the translated span. Consequently, if $i^{tgt}_{m}\notin s^{tgt}$, we prompt the LLM to find the span $i^{cor}_{m}$ in $s^{tgt}$ that semantically matches with the $i^{src}_{m}$. If $i^{cor}_{m}\in s^{tgt}$, $i^{tgt}_{m}$ will be rephrased as $i^{cor}_{m}$.
Stage 3: Sentence Rephrase
This stage focuses on the rephrasing of the translated sentence. If not all $i^{tgt}_{m}\in s^{tgt}$, we prompt the LLM to rephrase the $s^{tgt}$ based on the $s^{src}$ to ensure all $i^{tgt}_{m}$ are included.
Following Parekh et al. (2024), we evaluated the pipeline on the label projection benchmark WikiANN Pan et al. (2017), a multilingual IE dataset, and achieved average $99\%$ faithfulness in $10$ English-centric bilingual settings. The detailed results and settings are shown in Appendix A.1.
3.1.3 ParallelNER
Since alignment training requires including bilingual IE samples within a single prompt, the excessive concepts in RE and EE result in insufficient context length of LLMs. Thus, we construct NER parallel data for the phase. NER serves as a shared and foundational task in the IE framework, enabling entity identification through spotting abilities to support all IE tasks Lu et al. (2022).
[…]
3.2 Multilingual IE Training Phase
Following the previous phase, we conducted multilingual IE instruction tuning with $46$ IE datasets in English and Chinese. The detailed dataset statistics are shown in Appendix B. We follow the instruction and completion format introduced in Section 2.
4 Experimental Settings
AlignXIE enhances IE cross-lingual transfer by aligning the schema and extraction process. Our extensive evaluation of IE cross-lingual transfer can be divided into two main dimensions: external-language transfer (cross-lingual IE), and internal-language transfer (multilingual IE).
4.1 Metrics
We use span-based offset Micro-F1 to evaluate the methods. For NER, an entity is considered correct if the span and type match a golden annotation. For RE, a relation is considered correct if its type, subject entity, and object entity match a golden annotation. For ED, an event is valid if its trigger and type match a golden annotation. For EAE, given event type and trigger, an argument is valid if the span and its role match a golden annotation.
Model  Cross-Lingual Avgcross Avg
English Chinese German Spanish Dutch Russian Bengali Persian Hindi Korean Turkish
ChatGPT 37.20 18.80 37.10 34.70 35.70 27.40 23.30 25.90 27.30 30.00 31.90 30.37 29.94
GLiNER Zaratiana et al. (2024) 41.70 24.30 39.50 42.10 38.90 33.30 25.90 30.20 27.80 28.70 30.00 32.93 32.95
B2NER Yang et al. (2024) 54.80 45.40 36.60 46.00 43.00 33.90 - - - - - - -
[…]
Supervised Malmasi et al. (2022a) 62.70 53.10 64.60 58.70 62.60 59.70 39.70 52.30 47.80 55.80 46.80 54.22 54.89
Table 2: Results of cross-lingual evaluation in Multiconer22. Avgcross represents the average performance across $9$ unseen languages. indicates that the score of ChatGPT is reported by Lai et al., 2023. Supervised denotes the supervised method which is a strong baseline since the above four methods are under zero-shot setting.
4.2 Implementation Details
AlignXIE is fine-tuned based on Baichuan2-7B-Base Baichuan (2023). We utilize the LLaMA-Factory Zheng et al. (2024) for instruction tuning. We use LoRA Hu et al. (2022) for parameter-efficient fine-tuning and set the LoRA rank to $32$. The warmup ratio is set to $0.01$. The learning rates for these two phases are set to $3\times 10^{-4}$. We limit the sequence length to $4096$ and set the batch size to $256$. During the inference phase, we set the temperature to 00. All experiments are conducted on 8 x NVIDIA-A800 80G.
Dataset BERT YAYI-UIE IEPILE B2NER AlignXIE
ACE 2005 87.30 81.78 81.86 83.04 87.49
AnatEM 85.82 76.54 87.21 89.18 89.19
BC2GM 80.90 82.05 80.73 81.95 84.49
BC4CHEMD 86.72 88.46 90.45 88.96 89.57
BC5CDR 85.28 83.67 88.07 88.52 88.46
Broad Twitter 58.61 83.52 - 82.16 82.36
CoNLL 2003 92.40 96.77 92.49 92.56 94.69
FabNER 64.20 72.63 77.07 78.82 83.19
FindVehicle 87.13 98.47 98.49 97.89 99.47
GENIA 73.30 75.21 76.66 76.43 78.97
HarveyNER 82.26 69.57 67.70 73.67 73.91
MIT Movie 88.78 70.14 88.23 90.78 89.50
MIT Restaurant 81.02 79.38 79.85 83.71 81.95
MultiNERD 91.25 88.42 94.60 93.98 95.94
NCBI 80.20 87.29 85.26 84.83 85.49
OntoNotes 91.11 87.04 87.55 84.31 87.91
PolyglotNER 75.65 70.85 - 61.96 64.47
TweetNER7 56.49 66.99 - 66.26 67.98
WikiANN 70.60 72.63 - 85.07 84.69
wikiNeural 82.78 87.63 - 93.01 87.79
Avg 80.09 80.95 - 83.85 84.88
MSRA 94.95 95.97 87.99 92.22 96.01
ResumeNER 95.93 - 93.92 95.90 96.05
Avg 94.72 - 90.96 94.06 96.03
Table 3: Results of supervised evaluation on NER. Results of BERT on English benchmarks are from Wang et al. (2023a), and on Chinese benchmarks are from Li et al. (2020a).
4.3 Datasets
We followed the IEPILE Gui et al. (2024), YAYI-UIE Xiao et al. (2023), and B2NER Yang et al. (2024) to conduct a comprehensive evaluation . Overall, for English IE, we evaluate performance across $28$ benchmarks for NER, $10$ benchmarks for RE, $6$ benchmarks for ED, and $3$ benchmarks for EAE. For Chinese IE, we evaluate performance on $4$ benchmarks for NER, $5$ benchmarks for RE, $4$ benchmarks for ED, and $2$ benchmarks for EAE. For cross-lingual IE, we evaluate performance on the multilingual NER benchmark Multiconer22 Malmasi et al. (2022b).
We conduct supervised evaluations on various benchmarks: $23$ NER, $8$ RE, $3$ ED, and $3$ EAE benchmarks in English, as well as $2$ NER, $2$ RE, $2$ ED, and $2$ EAE benchmarks in Chinese. The detailed statistics and usage are presented in Appendix B. For the zero-shot evaluation, among NER benchmarks, we use $5$ English benchmarks from CrossNER Liu et al. (2021) and $2$ Chinese benchmarks, Weibo Peng and Dredze (2015) and Boson [^1] https://github.com/InsaneLife/. Among RE benchmarks, we adopt $2$ English benchmarks FewRel (Han et al., 2018), Wiki-ZSL (Chen and Li, 2021) and $3$ Chinese benchmarks, COAE2016222https://github.com/Sewens/COAE2016, IPRE (Wang et al., 2019), and SKE2020333https://aistudio.baidu.com/datasetdetail/177191. Among ED benchmarks, we use $3$ English benchmarks CrudeOil News Lee et al. (2021), RAMS Ebner et al. (2020), WikiEvents Li et al. (2021) and $2$ Chinese benchmarks FewFC (Zhou et al., 2021) and CCF law [^4] https://aistudio.baidu.com/projectdetail/4201483.
4.4 Baselines
We mainly use two types of baselines for a comprehensive comparison: Following InstructUIE Wang et al. (2023a), we first compare AlignXIE with BERT-Base Devlin (2018) methods. Additionally, we compare AlignXIE with the SoTA multilingual IE systems based on LLMs, including IEPILE Gui et al. (2024), YAYI-UIE Xiao et al. (2023), and B2NER Yang et al. (2024), which are based on natural language prompts. The three baselines exhibit varying degrees of unfairness when compared to AlignXIE . IEPILE and YAYI-UIE are fine-tuned based on the Baichuan2-13B, while B2NER uses Intern2LM-7B Cai et al. (2024).
For the cross-lingual evaluation, we compare with ChatGPT (evaluated by Lai et al., 2023), GLiNER Zaratiana et al. (2024), and B2NER. Furthermore, we further compare with the supervised baseline in Multiconer22 Malmasi et al. (2022b).
5 Results
5.1 Cross-Lingual Evaluation
To evaluate whether AlignXIE improves IE cross-lingual alignment, we conducted the cross-lingual evaluation on $9$ unseen languages of Multiconer22. We have also supplemented the evaluation in Chinese and English. Table 2 shows results. AlignXIE significantly outperforms ChatGPT and GLiNER in the cross-lingual setting, achieving a $30.17\%$ and $20.03\%$ improvement, respectively. For $4$ unseen languages used in B2NER that constitute more than $0.1\%$ of the general LLM pretraining corpus Touvron et al. (2023c), AlignXIE outperforms B2NER by $19.36\%$, with results comparable to B2NER in both Chinese and English, demonstrating a significant improvement in cross-lingual transfer despite lacking supervised training data as described in §4.4.
Dataset YAYI-UIE IEPILE AlignXIE
ADE corpus 84.14 83.73 84.45
CoNLL 2004 79.73 72.87 73.14
GIDS 72.36 74.71 76.19
kbp37 59.35 65.09 72.94
NYT 89.97 93.00 96.08
NYT11-HRL 57.53 53.19 56.79
SciERC 40.94 43.53 44.93
Semeval RE 61.02 58.47 64.79
Avg 68.13 68.07 71.16
CMeIE - 49.16 52.37
DuIE 2.0 81.19 75.61 82.85
Avg - 62.39 67.26
Table 4: Results of supervised evaluation on RE.
Surprisingly, despite not having any training data in $9$ unseen languages, AlignXIE has achieved performance comparable to the supervised baseline in Spanish and Turkish. The results demonstrate that AlignXIE exhibits excellent cross-lingual generalization for IE, which can be attributed to the enhancement of IE cross-lingual alignment.
Task Dataset BERT YAYI-UIE IEPILE AlignXIE
ED ACE 2005 72.50 65.00 72.46 73.57
CASIE 68.98 63.00 60.07 63.91
PHEE - 63.00 63.22 67.03
Avg - 63.67 65.25 68.17
DuEE 1.0 82.18 85.00 86.73 87.18
DuEE-Fin 84.53 82.50 83.54 85.13
Avg 83.36 83.75 85.14 86.16
EAE ACE 2005 59.90 62.71 63.90 69.95
CASIE 60.37 64.23 56.07 64.96
PHEE - 77.19 70.85 76.24
Avg - 68.04 63.61 70.67
DuEE 1.0 70.68 78.08 75.63 82.12
DuEE-Fin 75.73 70.02 79.34 81.09
Avg 73.21 74.05 77.49 81.61
Table 5: Results of supervised evaluation on EE. Results of BERT are from Xiao et al. (2023).
5.2 Supervised Evaluation
The results for NER, RE, EE (including ED and EAE) tasks are shown in Tables 3, 4, and 5 respectively. AlignXIE outperforms the multilingual SoTA baselines on most benchmarks for four tasks and ranks within the top-$2$ results across RE and ED benchmarks. In the English IE, AlignXIE has achieved significant average improvements of 3.03 and 2.92 F1 points on the RE and ED tasks, with an improvement of $7.85$ points on the kbp37 of RE. In the Chinese IE, AlignXIE has consistently achieved SoTA, with $4.12$ points average improvement over the SoTA baseline on the EAE task. Moreover, it can be observed that the model not only achieves significant improvements in NER but also exhibits even greater enhancements in RE, ED, and EAE. This indicates that the alignment phase enhances the multilingual NER capability through cross-lingual alignment, which improves the performance of all tasks by leveraging the foundational spotting capabilities of NER.
To further demonstrate that AlignXIE improves multilingual IE across all languages through IE cross-lingual alignment, we also compared it comprehensively with SoTA monolingual (English) IE systems, which are fine-tuning on LLMs including InstructUIE Wang et al. (2023a), UniversalNER Zhou et al. (2024), GoLLIE Sainz et al. (2024), KnowCoder Li et al. (2024), GLiNER Zaratiana et al. (2024), and GNER Ding et al. (2024). Appendix A.2 shows a detailed comparison with all baselines. Compared to $10$ SoTA methods in the supervised setting, it achieves top-$2$ results on $14$ (of $23$) English NER benchmarks, indicating substantial enhancements. Specifically, when compared to other code-based IE systems, such as KnowCoder and GoLLIE, our model still achieves a notable improvement. This observation underscores the effectiveness of cross-lingual alignment of AlignXIE . This indicates that our cross-lingual alignment not only avoids conflicts between different languages but also enhances overall multilingual IE performance by leveraging cross-language alignment.
Task Dataset GPT-4 YAYI-UIE IEPILE AlignXIE
NER AI - 52.40 - 65.04
Literature - 45.99 - 59.68
Music - 51.20 - 62.97
Politics - 51.82 - 63.65
Science - 50.53 - 62.29
Avg 58.49 50.39 55.55 62.73
RE FewRel 22.43 36.09 41.28 42.18
Wiki-ZSL 23.76 41.07 37.61 42.13
ED CrudeOil News 26.13 12.45 36.61 33.47
WikiEvents 5.24 10.97 9.12 12.59
RAMS 10.14 18.87 20.19 20.71
Avg - 38.02 37.14 42.26 46.47
Table 6: Results of zero-shot evaluation on English Benchmarks. Results of GPT-4 are from Gui et al. (2024)
Task Dataset GPT-4 YAYI-UIE IEPILE AlignXIE
NER Boson 48.15 49.25 55.77 59.58
Weibo 29.80 38.03 36.46 44.01
RE COAE2016 41.15 19.97 47.43 48.79
IPRE 18.15 22.97 29.76 32.43
SKE2020 56.77 70.80 72.50 72.91
ED CCF Law 42.12 12.87 63.53 68.90
FewEC 74.25 81.28 83.59 84.87
Avg - 44.34 42.17 55.58 58.78
Table 7: Results of zero-shot evaluation on Chinese Benchmarks. donates the result of Weibo using four types following the setups of the two baseline models, and the result for the eight types is $38.80$.
5.3 Zero-Shot Evaluation
Tables 6 and 7 show the zero-shot performance in English and Chinese, respectively. Since our method uses the data and setups of IEPILE, which differ from B2NER in data pruning, exclusion of dataset names, etc., the comparison with B2NER in zero-shot evaluation is unfair. These differences, as shown by Yang et al. 2024, significantly impact zero-shot performance, with the average gaps in English tasks reaching up to $9.3$ points.
AlignXIE achieved SoTA performance across all benchmarks in Chinese tasks, demonstrating substantial and consistent improvements in generalization ability. In English tasks, AlignXIE exhibited a noteworthy average increase of $4.24$ points on the NER task. Similarly, in Chinese tasks, AlignXIE surpassed existing SoTA models on all benchmarks, achieving an average improvement of $3.29$ points. Moreover, the trends demonstrated in both Chinese and English show that the most significant improvement occurs in NER. We hypothesize that the NER alignment training in the first phase is more challenging to generalize to other IE tasks in out-of-domain scenarios, which leaves the introduction of RE and EE alignment training to future work.
Dataset Text-Based Code-Based
ACE 2005 Walker et al. (2006) 84.93 86.07
AnatEM Pyysalo and Ananiadou (2014) 89.63 90.14
BC2GM Kocaman and Talby (2020) 82.53 83.17
BC5CDR Li et al. (2016) 90.03 90.17
CoNLL 2003 Sang and Meulder (2003) 94.67 94.93
WNUT 2017 Derczynski et al. (2017) 66.14 66.03
ResumeNER Zhang and Yang (2018) 94.31 95.37
MSRA Levow (2006) 92.97 94.03
Avg 86.90 87.49
Table 8: Ablation study on code-based method.
5.4 Ablation Study
We conduct comprehensive ablation experiments to investigate whether the code-based multilingual IE and IE cross-lingual alignment phase of AlignXIE contribute to the performance.
Code-Based Multilingual IE
We first evaluate the text-based Wang et al. (2023a) and code-based multilingual IE. Table 8 shows results. To ensure a fair comparison, we remove all class comments in our code-based instructions. We use $6$ English NER datasets, ACE 2005, AnatEM, BC2GM, BC5CDR, CoNLL 2003, WNUT 2017, and $2$ Chinese NER datasets, MSRA, ResumeNER to conduct instruction tuning on the Baichuan2-7B for supervised evaluation. The code-based method surpasses the text-based method by $0.59$ points, with increases of over $1.00$ points in $3$ datasets, especially the Chinese datasets MSRA and ResumeNER. The results show that the code can mitigate schema differences between languages, thereby enhancing the cross-lingual alignment to improve the performance of multilingual IE.
IE Alignment Training
To show whether the IE cross-lingual alignment phase contributes to the performance, we further conduct an ablation study by removing the cross-lingual alignment phase of AlignXIE and denoting it as AlignXIE (w/o align phase). Table 9 shows the average performance of supervised evaluation on $4$ tasks across $45$ benchmarks. The results indicate that, compared to AlignXIE (w/o align phase), AlignXIE achieves a substantial average improvement of $1.1$ points across four tasks. This suggests that the IE cross-lingual alignment phase enhances the overall performance of multilingual IE, demonstrating its contribution to facilitating cross-lingual transfer within the multilingual IE framework.
6 Related Work
Transfer Learning for cross-lingual IE
Huang et al. (2017) introduced a zero-shot learning method for EE, facilitating generalization across unseen languages. Building on zero-shot approaches, Huang et al. (2022) utilized generative language models for cross-lingual EAE. Hsu et al. (2023) enhanced cross-lingual RE through prompt-based learning, reducing reliance on large annotated datasets. Zubillaga et al. (2024) emphasized the significance of typology in cross-lingual EE, particularly for low-resource languages like Basque. In contrast, AlignXIE focuses on leveraging the spontaneous alignment of LLMs to facilitate cross-lingual IE transfer on all IE tasks.
Model NER RE ED EAE Avg
AlignXIE (w/o align phase) 85.06 68.67 74.12 74.00 75.46
AlignXIE 85.63 70.38 75.37 74.87 76.56
Table 9: Ablation study on IE alignment training.
LLMs for Multilingual IE
YAYI-UIE Xiao et al. (2023) proposes a chat-enhanced instruction tuning framework for UIE. IEPILE Gui et al. (2024) collects a comprehensive bilingual IE instruction corpus, and B2NER Yang et al. (2024) further designs a universal entity taxonomy. However, previous works overlook the mutual influence mechanisms between schemas in different languages and lack a unified representation.
Code-Based IE
Code-based IE aims to formulate the IE as a code generation task. CodeIE Li et al. (2023) first leverages the Code-LLM and recasts IE tasks into Python function completion tasks. Code4Struct Wang et al. (2023b) represents schema in Python classes for the EAE task. GoLLIE Sainz et al. (2024) incorporates guidelines within class comments. Besides, KnowCoder Li et al. (2024) utilizes class inheritance, class methods, and type hints to further enrich the schema representation. However, these works focus on the English IE task, and we are dedicated to exploring the significant role of code in multilingual IE.
7 Conclusion
In conclusion, this paper introduces AlignXIE , a code-based multilingual IE model that significantly enhances cross-lingual transfer through unified schema representation and an alignment training phase. AlignXIE utilizes Python classes to represent schemas uniformly across languages, ensuring semantic consistency in multilingual IE tasks. Additionally, our proposed cross-lingual alignment phase leverages high-quality IE parallel datasets ParallelNER, which are generated by our proposed LLM-based pipeline, to boost multilingual generalization. Comprehensive experiments demonstrate that AlignXIE achieves SoTA in Chinese and English IE under various settings, with remarkable cross-lingual capability.
Limitation
Extending multilingual IE training to languages beyond English and Chinese remains an exploration for future work. Meanwhile, the inclusion of additional languages would provide deeper insights into the cross-lingual alignment mechanisms within multilingual IE systems. Introducing parallel data from other IE tasks during the cross-lingual alignment phase also presents promising opportunities for further exploration. Furthermore, due to various limitations, we were unable to retrain the model or conduct a comprehensive evaluation using the B2NER framework. Addressing this issue will be a priority for future work.
Appendix A Experiments
A.1 Pipeline Evaluation
We evaluate our pipeline on the label projection benchmark WikiANN Pan et al. (2017), which is a multilingual NER dataset. We use $10$ languages, including Bengali (bn), German (de), Spanish (es), Persian (fa), Hindi (hi), Korean (ko), Dutch (nl), Russian (ru), Turkish (tr), and Chinese (zh), and randomly sample $1,000$ samples for each language. We compare our pipeline with CLaP Parekh et al. (2024),, Awesome-Align Dou and Neubig (2021), and EasyProject Chen et al. (2023). We mainly evaluate faithfulness following the setting of CLaP, measured as the percentage of instances where translated labels appear in the translated sentence, as accuracy evaluation requires costly native speaker rankings across methods, incurring high costs.
Table 10 shows the result. We achieve $99\%$ in faithfulness on average, outperforming the current SoTA Awesome-Align by $3\%$ points on average. We attribute the achievement to the latter two stages, which rephrase the span with minimal disturbance to the sentence. Besides, other baselines experience significant performance drops in certain languages, such as EasyProject in Russia. In contrast, our pipeline achieves scores exceeding $93\%$ across all languages, which strongly demonstrates the robustness and reliability of our pipeline. The results further substantiate the stability of our pipeline, demonstrating its capability to adapt effectively across different linguistic contexts.
Language Awesome-Align EasyProject CLaP Ours
bn 92 98 93 99
de 99 97 79 99
es 99 99 84 100
fa 96 99 72 100
hi 93 36 90 99
ko 96 93 64 99
nl 99 100 85 100
ru 97 99 66 93
tr 98 98 94 99
zh 92 92 60 99
Avg 96 91 79 99
Table 10: Results of label projection evaluation.
Dataset BERT InsturctUIE IEPILE YAYI-UIE GoLLIE UniversalNER GLiNER KnowCoder GNER B2NER AlignXIE
ACE 2005 87.30 79.94 81.86 81.78 89.10 86.69 82.80 86.10 - 83.04 87.49
AnatEM 85.82 85.82 87.21 76.54 - 88.65 88.90 86.40 90.24 89.18 89.19
BC2GM 80.90 80.69 80.73 82.05 - 82.42 83.70 82.00 83.18 81.95 84.49
BC4CHEMD 86.72 87.62 90.45 88.46 - 89.21 87.90 - 89.40 88.96 89.57
BC5CDR 85.28 89.02 88.07 83.67 91.90 89.34 88.70 89.30 90.27 88.52 88.46
Broad Twitter 58.61 80.27 83.52 - - 81.25 82.50 78.30 83.74 82.16 82.36
CoNLL 2003 92.40 91.53 92.49 96.77 92.90 93.30 92.60 94.10 93.60 92.56 94.69
FabNER 64.20 78.38 77.07 72.63 - 81.87 77.80 82.90 85.39 78.82 83.19
FindVehicle 87.13 87.56 98.49 98.47 - 98.30 95.70 99.40 98.62 97.89 99.47
GENIA 73.30 75.71 76.66 75.21 - 77.54 78.90 76.70 - 76.43 78.97
HarveyNER 80.20 74.49 67.70 69.57 - 74.21 - - 74.73 73.67 73.91
MIT Movie 88.78 89.58 88.23 70.14 - 90.17 87.90 89.70 90.23 90.78 90.24
MIT Restaurant 81.02 82.59 79.85 79.38 - 82.35 83.60 81.30 81.73 83.71 81.95
MultiNERD 91.25 90.26 94.60 88.42 - 93.73 93.80 96.10 94.30 93.98 95.94
NCBI 80.20 86.21 85.26 87.29 - 86.96 87.80 83.80 89.27 84.83 85.49
OntoNotes 91.11 88.64 87.04 87.55 - 89.91 - - 90.69 84.31 87.91
Polyglot 75.65 53.30 70.85 - - 65.70 61.50 - 67.52 61.96 64.47
TweetNER7 56.49 65.90 66.99 - - 65.80 51.40 - 66.87 66.26 67.98
WikiANN 70.60 64.47 72.63 - - 84.91 83.70 87.70 86.87 85.07 84.69
WikiNeural 82.78 88.27 87.63 - - 93.28 91.30 - 93.71 93.01 87.79
OntoNotes 5 84.28 - - - 83.40 - - 88.20 - - 88.35
DIANN 82.05 - - - 84.10 - - 94.70 - - 94.71
WNUT 2017 36.16 - - - 53.70 - - 66.40 - - 68.72
Table 11: Full comparison of supervised evaluation on NER. The results of BERT on DIANN, OntoNotes 5, and WNUT 2017 are from our implementation.
A.2 Full Comparison of Evaluation
We conduct a comprehensive comparison in English between AlignXIE and all current SoTA methods including monolingual IE methods and multilingual IE methods under the supervised setting in Tables 111213 and zero-shot setting in Table 14. In the NER task, we demonstrate top-$2$ results across $14$ (of $23$) benchmarks; in the RE task, we achieve top-$2$ results across $5$ (of $8$) benchmarks. Moreover, particularly in the ED and EAE tasks, we rank among the top-$2$ results across all benchmarks, further substantiating the significant efficacy of our method in multilingual IE.
Additionally, in the comparison of three benchmarks DIANN Zavala et al. (2018), Ontonotes 5 Hovy et al. (2006), and WNUT 2017 Derczynski et al. (2017), utilized by other English code-based IE work Sainz et al. (2024); Li et al. (2024), AlignXIE outperforms the other baselines, which further demonstrated that cross-lingual alignment can enhance monolingual IE.
In the zero-shot setting, we mainly compare our method with multilingual IE systems including IEPILE, and YAYI-UIE. Due to numerous different settings with B2NER, it is unfair to compare our method with B2NER. Our method achieves superior performance compared to other baseline approaches in the AI subset.
Appendix B Data Statistics
B.1 Dataset Statistics
In this work, we conduct evaluations on $63$ datasets, comprising $34$ datasets for the NER task, $15$ datasets for the RE task, $10$ datasets for the ED task, and $5$ datasets for the EAE task. To fairly compare with code-based methods Knowcoder and GoLLIE in demonstrating the effectiveness of our cross-language alignment, we also used PileNER Zhou et al. (2024) for training. The detailed statics are shown in Tables 15, 16, and 17, respectively.
B.2 ParallelNER
In this section, we mainly introduce the statistics of ParallelNER, which is constructed from two datasets WikiNeural Tedeschi et al. (2021) and CLUENER2020 Xu et al. (2020). The detailed statistics are shown in Table 18.
Appendix C Prompts
Description Generation
Using entity description generation as an example, we introduce the prompts we used including Description Initialization and Description Polish in Tables 19 and  20.
Dataset YAYI-UIE IEPILE InsturctUIE KnowCoder AlignXIE
ADE corpus 84.14 83.73 82.31 84.30 84.45
CoNLL 2004 79.73 72.87 78.48 73.30 73.14
GIDS 72.36 74.71 76.90 78.00 76.19
kbp37 59.35 65.09 36.14 73.20 72.94
NYT 89.97 93.00 91.00 93.70 96.08
NYT11-HRL 57.53 53.19 56.06 - 56.79
SciERC 40.94 43.53 37.40 40.00 44.93
Semeval RE 61.02 58.47 73.23 66.30 64.79
Table 12: Full comparison of supervised evaluation on RE.
Task Dataset BERT YAYI-UIE IEPILE KnowCoder AlignXIE
ED ACE 2005 72.50 65.00 72.46 74.2 73.57
CASIE 68.98 63.00 60.07 - 63.91
PHEE - 63.00 63.22 - 67.03
EAE ACE 2005 59.90 62.71 63.90 70.30 69.95
CASIE 60.37 64.23 56.07 - 64.96
PHEE - 77.19 70.85 - 76.24
Table 13: Full comparison of supervised evaluation on ED and EAE.
Dataset AI Literature Music Politics Science Avg
GPT4 - - - - - 58.49
YAYI-UIE 52.40 45.99 51.20 51.82 50.53 50.39
IEPILE - - - - - 55.55
B2NER 64.10 69.00 76.40 75.50 76.70 72.34
AlignXIE 65.04 59.68 62.97 63.65 62.29 62.73
InsturctUIE 48.40 48.80 54.40 49.90 49.40 50.18
GoLLIE 60.30 67.10 64.50 60.80 60.50 62.64
UniversalNER 53.50 59.40 65.00 60.80 61.10 59.96
KnowCoder 60.30 61.10 70.00 72.20 59.10 64.54
GliNER 60.60 68.40 69.50 74.80 69.40 68.54
GNER 63.10 68.20 75.70 69.40 69.90 69.26
Table 14: Full comparison of supervised evaluation on Zero-shot NER.
IE Parallel Data Construction Pipeline
[…]
#Dataset #Types #Major Domain #Train #Test #Train Num #Valid Num #Test Num
ACE2005 Walker et al. (2006) 7 News 7,299 964 1,060
AnatEM Pyysalo and Ananiadou (2014) 1 Biomedical 5,861 2,081 3,830
BC2GM Kocaman and Talby (2020) 1 Biomedical 12,500 2,500 5,000
BC4CHEMD Kocaman and Talby (2020) 1 Biomedical 30,488 30,468 26,204
BC5CDR Li et al. (2016) 2 Biomedical 4,560 4,581 4,797
Broad Twitter Derczynski et al. (2016) 3 Social Media 5334 2,000 2,001
CoNLL2003 Sang and Meulder (2003) 3 News 14,041 3,250 3,453
FabNER Kumar and Starly (2021) 12 Science 9,435 2,182 2,064
FindVehicle Guan (2022) 21 Traffic 21,547 20,777 20,769
GENIA Kim et al. (2003) 5 Biomedical 15,023 1,669 1,854
HarveyNER Chen et al. (2022) 4 Social Media 3,553 1,270 1,260
MIT Movie Liu et al. (2019) 12 Social Media 9,774 2,442 2,442
MIT Restaurant Liu et al. (2019) 8 Social Media 7,659 1,520 1,520
MultiNERD Tedeschi and Navigli (2022) 16 Wikipedia 134,144 10,000 10,000
NCBI Dogan et al. (2014) 1 Biomedical 5,432 923 940
OntoNotes Pradhan and Xue (2009) 18 General 107,032 14,110 10,838
PolygtNER Al-Rfou et al. (2014) 3 Wikipedia 393,941 - 10,000
TweetNER7 Ushio et al. (2022) 7 Social Media 1,325 7,111 576
WikiANN Pan et al. (2017) 3 Wikipedia 20,000 - 10,000
WikiNeural Tedeschi et al. (2021) 3 Wikipedia 92,720 11,590 11,597
DIANN Zavala et al. (2018) 1 Science 3,900 975 1,334
OntoNotes5 Hovy et al. (2006) 18 General 54994 7997 7782
WNUT 2017 Derczynski et al. (2017) 6 General 3,394 1,008 1,287
PileNER Zhou et al. (2024) 15,578 General 45,883 - -
CrossNER_AI Liu et al. (2021) 13 AI 100 350 431
CrossNER_literature Liu et al. (2021) 11 Literary 100 400 416
CrossNER_music Liu et al. (2021) 12 Musical 100 380 465
CrossNER_politics Liu et al. (2021) 8 Political 199 540 650
CrossNER_science Liu et al. (2021) 16 Scientific 200 450 543
MSRA Levow (2006) 3 News 46,364 4,500 3,442
ResumeNER Zhang and Yang (2018) 8 News 3,821 - 477
Weibo Peng and Dredze (2015) 8 Social media 1,350 - 270
Boson [^5] https://aistudio.baidu.com/datasetdetail/177191 6 News - - 191
MulticonerMalmasi et al. (2022b) 6 News - - 51,793
Table 15: Statistics of NER datasets.
#Dataset #Types #Major Domain #Train #Test #Train Num #Valid Num #Test Num
ADE corpus Gurulingappa et al. (2012) 5 Biomedical 3417 427 428
CoNLL 2004 Roth and Yih (2004) 5 News 922 231 288
GIDS Jat et al. (2018) 4 News 8526 1417 4307
NYT Riedel et al. (2010) 24 News 56,196 5000 5000
NYT11 Takanobu et al. (2018) 12 News 60765 - 362
kbp37 Zhang and Wang (2015) 18 News 15917 1724 3405
SciERC Luan et al. (2018a) 7 Scientific 1861 275 551
semeval RE Hendrickx et al. (2010) 10 Scientific 6507 1493 2717
FewRel Han et al. (2018) 100 Wikipedia - - 17291
Wiki-ZSL Chen and Li (2021) 83 Wikipedia - - 23113
DuIE2.0 Luan et al. (2018b) 48 News 171126 20652 -
CMeIE Guan et al. (2020) 53 Biomedical 14339 3585 -
COAE2016666https://github.com/Sewens/COAE2016 9 General - - 971
IPRE Wang et al. (2019) 35 General - - 3340
SKE2020777https://aistudio.baidu.com/datasetdetail/177191 49 News - - 3601
Table 16: Statistics of RE datasets. Since the test set for DuIE2.0 is not open-sourced, we use the valid set as our evaluation data.
#Dataset #Types #Major Domain #Train #Test #Train Num #Valid Num #Test Num
ACE05 Walker et al. (2006) 33 News 19216 901 676
CASIE Satyapanich et al. (2020) 5 Cybersecurity 3732 777 1492
PHEE Sun et al. (2022) 2 Biomedical 2897 960 968
CrudeOilNews Lee et al. (2021) 18 Oil News - - 356
RAMS Ebner et al. (2020) 106 News - - 887
WikiEvents Li et al. (2021) 31 Wikipedia - - 249
DuEE1.0 Li et al. (2020b) 65 News 11908 1492 -
DuEE-fin Han et al. (2022) 13 Finance 7015 1171 -
CCF_law [^8] https://aistudio.baidu.com/projectdetail/4201483 9 Law - - 971
FewFC Zhou et al. (2021) 5 Finance - - 2879
Table 17: Statistics of EE datasets. Since the test sets for DuEE1.0 and DuEE-fin are not open-sourced, we use the valid sets as our evaluation data.
#Source #Origin Language #Types #Instances #Train Num #Valid Num #Test Num #Total Num
WikiNeural En 3 149,011 92,720 11,590 11,597 115,907
CLUENER2020 Zh 10 19,720 10,000 1,343 1,345 12,688
Table 18: Statistics of ParallelNER datasets.
{CJK}
UTF8gkai
Prompt for Description Initialization.
# Writing Entity Descriptions
## Introduction
This guide provides a step-by-step process for writing a clear, concise, and accurate description of an entity type based on a provided list of examples. The objective is to generalize the shared characteristics of the examples without referencing any specific instance, giving a broad and comprehensive understanding of the entity type.
## Prerequisites
Before you begin, make sure you have the following:
- **Entity Type**: The name of the entity type that requires a description.
- **Entity List**: A set of examples representing this entity type.
- **Basic Description**: While not mandatory, familiarity with the general concept of the entity type could be beneficial.
## Step-by-Step Instructions
### Step 1: Begin with the Required Phrase
Each description should start with:
*"[Entity Type] refers to"**
This ensures consistency across all descriptions. Replace **[Entity Type]** with the actual type name.
### Step 2: Generalize the Shared Characteristics
- Review the **Entity List** to identify common traits among all examples.
- Avoid referring to specific examples directly. Generalize to cover the entire group.
- Example: If the list includes various vehicles (cars, trucks), the description should focus on common traits such as modes of transportation designed for movement.
### Step 3: Provide Comprehensive Coverage
- The description should encapsulate all critical aspects represented in the example list, accounting for any outliers or unusual cases.
- Example: If the list includes motorized vehicles and non-motorized bicycles, ensure the description covers both.
### Step 4: Output the Description
- After completing the description, present it without referencing explicit examples. It should summarize the entity type in a single, generalized statement.
- If you want to revise the description, output the modified description in the format.
## Conclusion
By following these steps, you will create an accurate, clear, and generalized description of an entity type. Start with the required phrase, focus on generalization, and keep the language simple yet precise.
## Input
Entity Type: {entity_type}
Entity Example List: {entity_example_list}
## Example Template for Output
Entity Type: {entity_type}
Entity Example List: {entity_example_list}
Entity Type Description: "{entity_type} refers to…"(in the language of the Entity list)
Table 19: Prompt for Description Initialization.
{CJK}
UTF8gkai
Prompt for Description Polish.
# Evaluating and Revising Entity Description
## Introduction
This guide provides a systematic approach to evaluate whether a given description accurately represents the characteristics of an entity type. If the description is accurate and complete, no revision is necessary. However, if inaccuracies or omissions exist, revisions are required to ensure clarity and consistency in classifying entities.
## Step-by-Step Instructions
### Step 1: Analyze the Entity Type Description
- Carefully review the **entity type description** provided.
- Example: For the entity type “Animal,” the description may include “living organisms that move, breathe, and consume organic matter.”
### Step 2: Analyze the Entity
- Review the specific entity’s characteristics, noting its unique features.
- Example: If the entity is “dog,” note traits like “mammal, four-legged, domesticated, etc.”
### Step 3: Evaluate the Description’s Accuracy and Completeness
- Compare the entity type description with the entity’s characteristics.
         - Does the description fully encompass the defining features of the entity?
         - Are any characteristics missing or misrepresented?
- Check for completeness:
         - Does the description cover all essential traits necessary for classification?
- Verify accuracy:
         - Are the described attributes factually correct?
### Step 4: Revise the Description (if necessary)
- If the description is incomplete or inaccurate, revise it to reflect the entity’s correct characteristics.
- Ensure the revised description is clear, precise, and free from ambiguities.
## Conclusion
Following these steps will ensure each entity’s description is both accurate and comprehensive. This process maintains clarity and consistency in classifying entities under their respective types.
## Input
Entity Type: {entity_type}
Entity Example List: {entity_example_list}
## Example Template for Output
Entity Type: {entity_type}
Entity Example List: {entity_example_list}
Entity Type Description: "{entity_type} refers to…"(in the language of the Entity list)
Table 20: Prompt for Description Polish.
{CJK}
UTF8gkai
Prompt for Translation.
Translate the sentence and spans from English to Chinese.
Please follow these guidelines:
1. Translate each span considering the context of the sentence.
2. Ensure the number of spans after translation matches the original number of spans.
3. When outputting spans, ensure only to output the translation of each span.
The following is a few examples:
[English]
"sentence": "The EU rejected Germany’s call for a boycott of British lamb."
"spans": ["EU"] [Chinese]
"sentence": "欧盟拒绝德国呼吁抵制英国羊肉。"
"spans": ["欧盟"]
[English]
"sentence": "FM involves 2 - 4.7% of the general population."
"spans": [] [Chinese]
"sentence": "欧盟拒绝德国呼吁抵制英国羊肉。"
"spans": ["欧盟"]
[English]
"sentence": "4000 guests from home and abroad attended the opening ceremony."
"spans": ["home", "abroad"] [Chinese]
"sentence": "4000名来自国内和国外的嘉宾出席了开幕式。"
"spans": ["国内", "国外"]
Please translate the following sentence and spans:
[English]
"sentence": "{src_sentence}"
"spans": [{src_spans}] [Chinese]
Table 21: Prompt for Translation.
{CJK}
UTF8gkai
Prompt for Span Rephrase.
Please find the Chinese span corresponding to the English span in the Chinese sentence.
Please follow these guidelines:
1. Only find the span in the Chinese sentence that corresponds to the English span.
2. Ensure that the Chinese span must be semantically consistent with the English span.
The following is an example:
[English]
"sentence": "Siemens invested 800 million US dollars to complete the electric power plant project."
"spansZQ": ["US"] [Chinese]
"sentence": "西门子投资了8亿美元完成了电力厂项目。"
"spans": ["美"]
Please find the corresponding span in the Chinese sentence:
[English]
"sentence": "{src_sentence}"
"spans": [{src_span}] [Chinese]
"sentence": "{tgt_sentence}"
"spans":
Table 22: Prompt for Span Rephrase.
Prompt for Sentence Rephrase.
Please translate the following sentence from English to Chinese.
Please follow these guidelines:
1. Ensure that the translation includes the following spans: [{tgt_lang_spans}].
2. If the target sentence is semantically inconsistent with the source sentence, return "modification failure".
[English]
"sentence": "{src_sentence}"
[Chinese]
"sentence":
Table 23: Prompt for Sentence Rephrase.
Appendix D Examples
We present the examples of instruction-tuning data for the IE cross-lingual alignment phase and the multi-lingual IE training phase in Figure 2 and Figure 3, respectively.
Figure 2: An example of instruction-tuning data for the IE cross-lingual alignment phase.
Figure 3: An example of instruction-tuning data for the multilingual IE training phase.